December 20, 1995

Dr. Dinesh P. Mehta The University of Tennessee Space Institute Computer Science Department Mail Stop 21 Tullahoma, Tennessee 37388-8897

Dear Dr. Mehta,

Thank you very much for your interest in Concurrency: Practice and Experience. We enclose two referee reports and one marked up paper on, “GRIDMAP: A System for the Optimal Mapping of Grids onto Processors” (C269).

Both referees agree that significant changes are needed to focus on new results. We cannot publish your paper in its current form, but would be pleased to receive a revised version. Please enclose a memo describing changes, and your response to the referees. Also, resubmit your paper double-spaced with the tables and figures placed at the end of the text. Note in the text where each table and figure is to be placed.

Yours Sincerely,

Geoffrey C. Fox

Professor of Computer Science and Physics

Paul Messina

Director, Center for Advanced Computing Research

GCF/tc

Enclosure: Referee reports